OpenStudy (anonymous):

The roots of the equation 9x^2+3x-4=0 are 1) imaginary 2) real rational and equal 3) real rational and unequal 4) real irrational and unequal

OpenStudy (anonymous):

What do we get from the Quadratic formula? \[\frac {-3\pm \sqrt{9+144}}{18}=\]

OpenStudy (anonymous):

check \[b^2-4ac=9^2-4\times 9\times (-4)=225=15^2\] since the discriminant is a perfect square, you have two rational roots